how do you solve this
5(x-3)+4x-1=x+3(x-3)
: how do you solve this
5(x-3)+4x-1=x+3(x-3)

Answer by checkley77(3848)

5(x-3)+4x-1=x+3(x-3)
5x-15+4x-1=x+3x-9
9x-4x=-9+16
5x=7
x=7/5
x=7/5 ans.